# appendix,by

### v1

@@ 3.2 Amalgamation
If the amalgamation problem $M\leftarrow K\to L$ has a solution, then this clearly descends to the residue difference field level, so we prove the converse. Let $N_{0}$ be a valued difference field with one of the additional structures mentioned above, whose residue difference field $k(N_{0})$ solves the amalgamation problem $k(M)\leftarrow k(K)\to k(L)$. Let $N$ be an existentially closed, existentially $\lvert M\rvert^{+}-$saturated extension of $N_{0}$, so that $k(M)$ and $k(L)$ embed into $k(N)$ over $k(K)$. By Corollary˜2.22 and Corollary˜3.11, assumption (1) of Proposition˜3.27 is satisfied. Moreover, as $\Gamma(N)$ is existentially closed and existentially $\lvert\Gamma(M)\rvert^{+}-$saturated, it follows from [DM24, Theorem 5.10] that assumption (2) is satisfied as well, while assumption (3) is satisfied by hypothesis. Hence $M$ embeds into $N$ over $K$, solving the amalgamation problem.
For the “in particular” clause, the first part is clear. The fact that if the residue automorphism extends uniquely to the algebraic closure then $k(K)$ is an amalgamation base for difference fields follows from [Mac97, Lemma 8]. Conversely, if $(k(K),\sigma)$ is an amalgamation base and $\sigma_{1}$, $\sigma_{2}$ are two extensions to the algebraic closure, then any solution of the amalgamation problem $(k(K)^{\mathrm{alg}},\sigma_{1})\leftarrow(k(K),\sigma)\to(k(K)^{\mathrm{alg}},\sigma_{2})$ gives an isomorphism between $\sigma_{1}$ and $\sigma_{2}$. ∎
Solvable amalgamation problems appear in the difference algebra literature under the name of compatible extensions. See e.g. [Coh65, Chapter 7]. See [CH99, Lemmas 2.8 and 2.9(4)] for a condition ensuring that a field automorphism extends uniquely to the algebraic closure.
@@
If $(k,\sigma)$ and $(\Gamma,\sigma)$ are both existentially closed, then the Hahn field $k((\Gamma))$, with $\sigma$ and $\operatorname{ac}$ as in ˜1.6, is $\sigma$-henselian by [DO15, Corollary 5.6(2)]. By ˜3.29, it is existentially closed.
@@ 4 The tree property of the second kind
In this section we prove that, for $\mathcal{L}$ any of the languages considered above (see ˜3.22), the positive theory of $\mathcal{L}$-valued difference fields of residual characteristic zero is $\mathsf{NTP}_{2}$. We do this by adapting the proof of [CH14, Theorem 4.1]. This will require generalising certain statements about indiscernible arrays to the positive setting, which we do in the first two subsections. Along the way, we prove the Array Modelling Theorem for arbitrary positive theories, removing the thickness assumption used in [Kam24], and deduce submultiplicativity of burden from it. We also provide an example showing that Lemma 2.3 in [Che14] is false (this, however, does not affect the correctness of any other results of [Che14]).
Henceforth we assume familiarity with positive logic. We refer the reader to [BP07] for the fundamentals; condensed accounts are available for instance in [DK22, DGK23, DM24]. Terminology-wise, let us just mention that we say positive theory for h-inductive theory, and that we talk of positively existentially closed (pec) models, while “existentially closed” is reserved for the classical notion. We refer to [DGK23, Definitions 4.1 and 4.5] for the definitions of $\mathsf{IP}$ and $\mathsf{TP_{2}}$ in the positive setting. From now on, unless otherwise stated, we work in a monster model $\mathfrak{U}$ of a positive theory with the Joint Continuation Property (JCP). By a formula we will mean a positive formula, and by a type we will mean a (partial) positive type, we write $\equiv_{A}$ for having the same positive type over $A$, if we call a sequence indiscernible we mean with respect to positive formulas, etc. Similarly, every mention of $\mathsf{NTP}_{2}$ and similar properties refers to their versions for positive logic.
@@
Let $\mathcal{L}$ be a language and $I$, $J$ two $\mathcal{L}$-structures indexing some tuples $(a_{i})_{i\in I}$ and $(b_{j})_{j\in J}$.
@@
    The $\mathrm{EM}_{\mathcal{L}}$-type $\operatorname{EM}_{\mathcal{L}}((a_{i})_{i\in I}/A)$ of $(a_{i})_{i\in I}$ over $A$ is the set of formulas $\varphi(x_{\bar{\eta}})$ over $A$, as $\bar{\eta}$ varies among finite tuples in $I$, such that, for every tuple $\bar{\xi}$ from $I$ with the same quantifier-free $\mathcal{L}$-type as $\bar{\eta}$, we have $\vDash\varphi(a_{\bar{\xi}})$.
@@
    We say that $(a_{i})_{i\in I}$ is $\mathcal{L}$-indiscernible over $A$ iff
    $\operatorname{EM}_{\mathcal{L}}((a_{i})_{i\in I}/A)=\operatorname{tp}((a_{i})_{i\in I}/A).$
@@
    We say that $(b_{j})_{j\in J}$ is $\mathrm{EM}_{\mathcal{L}}$-based on $(a_{i})_{i\in I}$ over $A$ iff, for all $\varphi(x_{\bar{\eta}})\in\operatorname{EM}_{\mathcal{L}}((a_{i})_{i\in I}/A)$ and all tuples $\bar{\xi}$ from $J$ with the same quantifier-free $\mathcal{L}$-type as $\bar{\eta}$, we have $\varphi(b_{\bar{\xi}})\in\operatorname{EM}_{\mathcal{L}}((b_{i})_{i\in I}/A)$.
@@
    We say that $(b_{j})_{j\in J}$ is based on $(a_{i})_{i\in I}$ over $A$ iff, for all tuples $\bar{\xi}\subseteq J$ there is $\bar{\eta}\subseteq I$ such that $b_{\bar{\xi}}\equiv_{A}a_{\bar{\eta}}$.
@@
    $\mathrm{EM}_{\mathcal{L}}$-basedness is most significant (and will only be used below) when $I$ and $J$ realise the same finitary quantifier-free $\mathcal{L}$-types, in which case it is a transitive notion [Kam24, Remark 3.8]. Note that, if $I=J$, then being $\mathrm{EM}_{\mathcal{L}}$-based on” is simply inclusion of $\mathrm{EM}_{\mathcal{L}}$-types.
@@
    If $(b_{j})_{j\in J}$ is based on $(a_{i})_{i\in I}$ over $A$, then it is also $\mathrm{EM}_{\mathcal{L}}$-based on it over $A$, but the converse is false in general [Kam24, Proposition 3.10 and Example 3.11].
@@
When index sets are linear orders, we will write $\mathrm{EM}_{<}$ instead of $\mathrm{EM}_{\{<\}}$. In the same setting, the following fact is standard, see e.g. [Ben03a, Lemma 1.2].
@@
Let $A$ be a parameter set, $\kappa$ a cardinal, and let $\lambda\coloneqq\beth((2^{\lvert T\rvert+\lvert A\rvert+\kappa})^{+})$. Then for every sequence $(a_{i})_{i<\lambda}$ of compatible $\kappa$-tuples there is a sequence $(b_{i})_{i<\omega}$ that is $A$-indiscernible and based on $(a_{i})_{i<\lambda}$ over $A$.
@@
Fact 4.4 ([DK22, Corollary 5.4]).
Let $A$ be a parameter set. Then for every sequence $(a_{i})_{i<\omega}$ of compatible tuples there is a sequence $(b_{i})_{i<\omega}$ that is $A$-indiscernible and $\mathrm{EM}_{<}$-based on $(a_{i})_{i<\lambda}$ over $A$.
@@ 4.1 Array Modelling and submultiplicativity of burden
The first objective of this subsection is to generalise [Kam24, Theorems 1.3 and 1.4] to arbitrary positive theories, dropping the thickness assumption. We follow notation from there.
Recall that $\mathcal{L}_{\mathrm{ar}}$ is the language $\{<_{1},<_{2}\}$, interpreted in a subset $Z$ of a product of linear orders $I\times J$ as $(i,j)<_{1}(i^{\prime},j^{\prime})$ iff $i<i^{\prime}$ and $(i,j)<_{2}(i^{\prime},j^{\prime})$ iff $i=i^{\prime}$ and $j<j^{\prime}$. That is, $_{1}$ is the order on rows and $_{2}$ is the order internal to rows. For $\lambda$ an ordinal, $\mathcal{L}_{\lambda}$ denotes the expansion of $\mathcal{L}_{\mathrm{ar}}$ by predicates $P_{\alpha}$, for $\alpha<\lambda$, which we will interpret in a subset $Z$ of $\lambda\times\mu$ as $Z\cap(\{\alpha\}\times\mu)$.
If $Z$ is as above, we call $(a_{(ij)})_{(i,j)\in Z}$ an array. In order to keep the notation concise, we will also simply write $a_{Z}$, and extend this usage to restrictions, e.g. if $Z^{\prime}\subseteq Z$, we will write $a_{Z^{\prime}}$ with the obvious meaning. Similarly for e.g. $a_{<\alpha,\geq\beta}$. We will use $\bar{}$ to indicate rows, e.g. $\bar{a}_{\alpha}$ typically will mean the $\alpha$-th row of $a_{Z}$, and we similarly write e.g. $\bar{a}_{<\alpha}$ for the rows of $a_{Z}$ up to the $\alpha$-th, etc.
@@
    An array is $\mathcal{L}_{\lambda}$-indiscernible exactly when it is mutually indiscernible, i.e. each of its rows is indiscernible over the union of all other rows.
@@
    An array indexed on $I\times J$ is $\mathcal{L}_{\mathrm{ar}}$-indiscernible exactly when each of its rows is indiscernible over the union of all other rows and, furthermore, its rows form an indiscernible sequence.
We use strongly indiscernible as a synonym for $\mathcal{L}_{\mathrm{ar}}$-indiscernible”; note that [Kam24] uses the terminology array-indiscernible instead.
@@
Let $I$ be an $\mathcal{L}$-structure, $J,L$ be $\mathcal{L}^{\prime}$-structures, and $E$ a set of parameters. If $c_{L}$ is based on $b_{J}$ over $a_{I}E$ and $a_{I}$ is $\mathcal{L}$-indiscernible over $b_{J}E$, then $a_{I}$ is also $\mathcal{L}$-indiscernible over $c_{L}E$.
@@
By naming parameters, we may assume $E=\emptyset$. Consider any finite tuples $\bar{\eta},\bar{\nu}\subseteq I$ with $\operatorname{qftp}_{\mathcal{L}}(\bar{\eta})=\operatorname{qftp}_{\mathcal{L}}(\bar{\nu})$, as well as a finite tuple $\bar{\ell}\subseteq L$. As $c_{L}$ is based on $b_{J}$ over $a_{I}$, there is a tuple $\bar{\jmath}\subseteq J$ with $b_{\bar{\jmath}}\equiv_{a_{I}}c_{\bar{\ell}}$. As $a_{I}$ is $\mathcal{L}$-indiscernible over $b_{J}$, we have that $a_{\bar{\nu}}\equiv_{b_{\bar{\jmath}}}a_{\bar{\eta}}$. It follows that $a_{\bar{\nu}}\equiv_{c_{\bar{\ell}}}a_{\bar{\eta}}$, and, by compactness, that $a_{\bar{\nu}}\equiv_{c_{L}}a_{\bar{\eta}}$. ∎
For an ordinal $\lambda$ and $I\subseteq\lambda$, by $\operatorname{ot}(I)$ we will mean the order type of $I$ with the order induced from $\lambda$. For $Z\subseteq\lambda\times\mu$ we write $Z_{I}\coloneqq Z\cap(I\times\mu)$ and consider it as an $\mathcal{L}_{\operatorname{ot}(I)}$-structure in the obvious way.
By inspecting atomic $\mathcal{L}_{\lambda}$-formulas, we easily see that if $\lambda=I\dot{\cup}J$, $\bar{\nu},\bar{\nu}^{\prime}\subseteq Z_{I}$ and $\bar{\xi},\bar{\xi}^{\prime}\subseteq Z_{J}$, then $\operatorname{qftp}_{\mathcal{L}_{\lambda}}(\nu,\xi)=\operatorname{qftp}_{\mathcal{L}_{\lambda}}(\nu^{\prime},\xi^{\prime})$ if and only if $\operatorname{qftp}_{\mathcal{L}_{\operatorname{ot}(I)}}(\nu)=\operatorname{qftp}_{\mathcal{L}_{\operatorname{ot}(I)}}(\nu^{\prime})$ and $\operatorname{qftp}_{\mathcal{L}_{\operatorname{ot}(J)}}(\xi)=\operatorname{qftp}_{\mathcal{L}_{\operatorname{ot}(J)}}(\xi^{\prime})$.
@@
Let $\lambda=I\dot{\cup}J$ and $Z,Z^{\prime}\subseteq\lambda\times\mu$ with $Z_{J}=Z^{\prime}_{J}$. Suppose $b_{Z^{\prime}_{I}}$ is $\mathrm{EM}_{\mathcal{L}_{\operatorname{ot}(I)}}$-based on $a_{Z_{I}}$ over $a_{Z_{J}}E$. Let
@@
Then $c_{Z^{\prime}}$ is $\mathrm{EM}_{\mathcal{L}_{\lambda}}$-based on $a_{Z}$ over $E$.
@@
We may assume that $E=\emptyset$. Let $\bar{\nu}\subseteq Z_{I}$, $\bar{\xi}\subseteq Z_{J}$, and $\varphi$ be such that $\varphi(x_{\bar{\nu}},x_{\bar{\xi}})\in\operatorname{EM}_{\mathcal{L}_{\lambda}}(a_{Z})$. We have to prove that $\vDash\varphi(c_{\bar{\nu}},c_{\bar{\xi}})$.
For every $\bar{\nu}^{\prime}$ with $\operatorname{qftp}_{\mathcal{L}_{\operatorname{ot}(I)}}(\bar{\nu}^{\prime})=\operatorname{qftp}_{\mathcal{L}_{\operatorname{ot}(I)}}(\bar{\nu})$ we have that $\operatorname{qftp}_{\mathcal{L}_{\operatorname{ot}(I)}}(\bar{\nu}^{\prime},\bar{\xi})=\operatorname{qftp}_{\mathcal{L}_{\operatorname{ot}(I)}}(\bar{\nu},\bar{\xi})$, so $\vDash\varphi(a_{\bar{\nu}^{\prime}},a_{\bar{\xi}})$. Hence $\varphi(x_{\bar{\nu}^{\prime}},a_{\bar{\xi}})\in\operatorname{EM}_{\mathcal{L}_{\operatorname{ot}(I)}}(a_{Z_{I}}/a_{Z_{J}})$ and so $\vDash\varphi(b_{\bar{\nu}},a_{\bar{\xi}})$ by the assumption on $b_{Z^{\prime}_{I}}$. This means that $\vDash\varphi(c_{\bar{\nu}},c_{\bar{\xi}})$, as required. ∎
@@
Let $\lambda=I\dot{\cup}J$ and $Z\subseteq\lambda\times\mu$. Suppose $a_{Z}$ is such that $a_{Z_{J}}$ is $\mathcal{L}_{\operatorname{ot}(J)}$-indiscernible over $a_{Z_{I}}E$ and $a_{Z_{I}}$ is $\mathcal{L}_{\operatorname{ot}(I)}$-indiscernible over $a_{Z_{J}}E$. Then $a_{Z}$ is $\mathcal{L}_{\lambda}$-indiscernible over $E$.
@@
We may assume $E=\emptyset$. Let $\bar{\nu},\bar{\nu}^{\prime}\subseteq Z_{I}$ and $\bar{\xi},\bar{\xi}^{\prime}\subseteq Z_{J}$ be such that $\operatorname{qftp}_{\mathcal{L}_{\lambda}}(\bar{\nu},\bar{\xi})=\operatorname{qftp}_{\mathcal{L}_{\lambda}}(\bar{\nu}^{\prime},\bar{\xi}^{\prime})$. It follows easily from the assumptions that $a_{\bar{\nu},\bar{\xi}}\equiv a_{\bar{\nu},\bar{\xi}^{\prime}}\equiv a_{\bar{\nu}^{\prime},\bar{\xi}^{\prime}}$. ∎
@@
Let $A=(a_{ij})_{i<\lambda,j<\omega}$ be an array with each row consisting of compatible tuples, and $E$ a set of parameters. Then there is an array $(b_{ij})_{i<\lambda,j<\omega}$ which is $\mathcal{L}_{\lambda}$-indiscernible over $E$ and $\mathrm{EM}_{\mathcal{L}_{\lambda}}$-based on $A$ over $E$.

### v2

@@ 3.2 Amalgamation
If the amalgamation problem $M\leftarrow K\to L$ has a solution, then this clearly descends to the residue difference field level, so we prove the converse. Let $N_{0}$ be a valued difference field with one of the additional structures mentioned above, whose residue difference field $k(N_{0})$ solves the amalgamation problem $k(M)\leftarrow k(K)\to k(L)$. Let $N$ be an existentially closed, existentially $\lvert M\rvert^{+}-$saturated extension of $N_{0}$, so that $k(M)$ and $k(L)$ embed into $k(N)$ over $k(K)$. By Corollary˜2.22 and Corollary˜3.11, assumption (1) of Proposition˜3.27 is satisfied. Moreover, as $\Gamma(N)$ is existentially closed and existentially $\lvert\Gamma(M)\rvert^{+}-$saturated, it follows from [DM, Theorem 5.10] that assumption (2) is satisfied as well, while assumption (3) is satisfied by choice of $N_{0}$. Hence $M$ embeds into $N$ over $K$, solving the amalgamation problem.
For the “in particular” clause, the first part is clear. The fact that if the residue automorphism extends uniquely to the algebraic closure then $k(K)$ is an amalgamation base for difference fields follows from [Ma, Lemma 8]. Conversely, if $(k(K),\sigma)$ is an amalgamation base and $\sigma_{1}$, $\sigma_{2}$ are two extensions to the algebraic closure, then any solution of the amalgamation problem $(k(K)^{\mathrm{alg}},\sigma_{1})\leftarrow(k(K),\sigma)\to(k(K)^{\mathrm{alg}},\sigma_{2})$ induces an isomorphism between $(k(K)^{\mathrm{alg}},\sigma_{1})$ and $(k(K)^{\mathrm{alg}},\sigma_{2})$. ∎
Solvable amalgamation problems appear in the difference algebra literature under the name of compatible extensions. See e.g. [cohnDifferenceAlgebra1965, Chapter 7]. See [CHr, Lemmas 2.8 and 2.9(4)] for a condition ensuring that a field automorphism extends uniquely to the algebraic closure.
@@
If $(k,\sigma)$ and $(\Gamma,\sigma)$ are both existentially closed and $\operatorname{char}(k)=0$, then the Hahn field $k((\Gamma))$, with $\sigma$ and $\operatorname{ac}$ as in ˜1.6, is $\sigma$-henselian by [DO, Corollary 5.6(2)]. By ˜3.29, it is existentially closed.
@@ 4 The tree property of the second kind
In this section we prove that, for $\mathcal{L}$ any of the languages considered above (see ˜3.22), the positive theory of $\mathcal{L}$-valued difference fields of residual characteristic zero is $\mathsf{NTP}_{2}$. We do this by adapting the proof of [CH, Theorem 4.1]. This will require generalising certain statements about indiscernible arrays to the positive setting, which we do in the first two subsections. Along the way, we prove the Array Modelling Theorem for arbitrary positive theories, removing the thickness assumption used in [Ka], and deduce submultiplicativity of burden from it. We also provide an example showing that Lemma 2.3 in [Chernikov] is false (this, however, does not affect the correctness of any other results of [Chernikov]).
Henceforth we assume familiarity with positive logic. We refer the reader to [BYP] for the fundamentals; condensed accounts are available for instance in [DK, DGK, DM]. Terminology-wise, let us just mention that we say positive theory for h-inductive theory, and that we talk of positively existentially closed (pec) models, while “existentially closed” is reserved for the classical notion. We refer to [DGK, Definitions 4.1 and 4.5] for the definitions of $\mathsf{IP}$ and $\mathsf{TP_{2}}$ in the positive setting. From now on, unless otherwise stated, we work in a monster model $\mathfrak{U}$ of a positive theory with the Joint Continuation Property (JCP). By a formula we will mean a positive formula, and by a type we will mean a (partial) positive type, we write $\equiv_{A}$ for having the same positive type over $A$, if we call a sequence indiscernible we mean with respect to positive formulas, etc. Similarly, every mention of $\mathsf{NTP}_{2}$ and similar properties refers to their versions for positive logic.
@@
Let $\mathfrak{L}$ be a language and $I$, $J$ two $\mathfrak{L}$-structures indexing some tuples $(a_{i})_{i\in I}$ and $(b_{j})_{j\in J}$.
@@
    The $\mathrm{EM}_{\mathfrak{L}}$-type $\operatorname{EM}_{\mathfrak{L}}((a_{i})_{i\in I}/A)$ of $(a_{i})_{i\in I}$ over $A$ is the set of formulas $\varphi(x_{\bar{\eta}})$ over $A$, as $\bar{\eta}$ varies among finite tuples in $I$, such that, for every tuple $\bar{\xi}$ from $I$ with the same quantifier-free $\mathfrak{L}$-type as $\bar{\eta}$, we have $\vDash\varphi(a_{\bar{\xi}})$.
@@
    We say that $(a_{i})_{i\in I}$ is $\mathfrak{L}$-indiscernible over $A$ iff
    $\operatorname{EM}_{\mathfrak{L}}((a_{i})_{i\in I}/A)=\operatorname{tp}((a_{i})_{i\in I}/A).$
@@
    We say that $(b_{j})_{j\in J}$ is $\mathrm{EM}_{\mathfrak{L}}$-based on $(a_{i})_{i\in I}$ over $A$ iff, for all $\varphi(x_{\bar{\eta}})\in\operatorname{EM}_{\mathfrak{L}}((a_{i})_{i\in I}/A)$ and all tuples $\bar{\xi}$ from $J$ with the same quantifier-free $\mathfrak{L}$-type as $\bar{\eta}$, we have $\varphi(b_{\bar{\xi}})\in\operatorname{EM}_{\mathfrak{L}}((b_{i})_{i\in I}/A)$.
@@
    We say that $(b_{j})_{j\in J}$ is based on $(a_{i})_{i\in I}$ over $A$ iff, for all tuples $\bar{\xi}\subseteq J$ there is $\bar{\eta}\subseteq I$, with the same quantifier-free $\mathfrak{L}$-type as $\bar{\xi}$, such that $b_{\bar{\xi}}\equiv_{A}a_{\bar{\eta}}$.
@@
    $\mathrm{EM}_{\mathfrak{L}}$-basedness is most significant (and will only be used below) when $I$ and $J$ realise the same finitary quantifier-free $\mathfrak{L}$-types, in which case it is a transitive notion [Ka, Remark 3.8]. Note that, if $I=J$, then being $\mathrm{EM}_{\mathfrak{L}}$-based on” is simply inclusion of $\mathrm{EM}_{\mathfrak{L}}$-types.
@@
    If $(b_{j})_{j\in J}$ is based on $(a_{i})_{i\in I}$ over $A$, then it is also $\mathrm{EM}_{\mathfrak{L}}$-based on it over $A$, but the converse is false in general [Ka, Proposition 3.10 and Example 3.11].
@@
When index sets are linear orders, we will write $\mathrm{EM}_{<}$ instead of $\mathrm{EM}_{\{<\}}$. In the same setting, the following fact is standard, see e.g. [ben-yaacovSimplicityCompactAbstract2003, Lemma 1.2].
@@
Let $A$ be a parameter set, $\kappa$ a cardinal, and let [^15] For the sake of readability, we will use the notation $\beth(\alpha)$ instead of the more commonly used $\beth_{\alpha}$. $\lambda\coloneqq\beth((2^{\lvert T\rvert+\lvert A\rvert+\kappa})^{+})$. Then for every sequence $(a_{i})_{i<\lambda}$ of compatible $\kappa$-tuples there is a sequence $(b_{i})_{i<\omega}$ that is $A$-indiscernible and based on $(a_{i})_{i<\lambda}$ over $A$.
@@
Fact 4.4 ([DK, Corollary 5.4]).
Let $A$ be a parameter set. Then for every sequence $(a_{i})_{i<\omega}$ of compatible tuples there is a sequence $(b_{i})_{i<\omega}$ that is $A$-indiscernible and $\mathrm{EM}_{<}$-based on $(a_{i})_{i<\omega}$ over $A$.
@@ 4.1 Array Modelling and submultiplicativity of burden
The first objective of this subsection is to generalise [Ka, Theorems 1.3 and 1.4] to arbitrary positive theories, dropping the thickness assumption. We follow notation from there.
Recall that $\mathfrak{L}_{\mathrm{ar}}$ is the language $\{<_{1},<_{2}\}$, interpreted in a subset $Z$ of a product of linear orders $I\times J$ as $(i,j)<_{1}(i^{\prime},j^{\prime})$ iff $i<i^{\prime}$ and $(i,j)<_{2}(i^{\prime},j^{\prime})$ iff $i=i^{\prime}$ and $j<j^{\prime}$. That is, $_{1}$ is the order on rows and $_{2}$ is the order internal to rows. For $\lambda$ an ordinal, $\mathfrak{L}_{\lambda}$ denotes the expansion of $\mathfrak{L}_{\mathrm{ar}}$ by predicates $P_{\alpha}$, for $\alpha<\lambda$, which we will interpret in a subset $Z$ of $\lambda\times\mu$ as $Z\cap(\{\alpha\}\times\mu)$.
If $Z$ is as above, we call $(a_{(ij)})_{(i,j)\in Z}$ an array. In order to keep the notation concise, we will also simply write $a_{Z}$, and extend this usage to restrictions, e.g. if $Z^{\prime}\subseteq Z$ we will write $a_{Z^{\prime}}$ with the obvious meaning. Similarly for e.g. $a_{<\alpha,\geq\beta}$. We will use $\bar{}$ to indicate rows, e.g. $\bar{a}_{\alpha}$ typically will mean the $\alpha$-th row of $a_{Z}$, and we similarly write e.g. $\bar{a}_{<\alpha}$ for the sequence whose elements are the rows of $a_{Z}$ up to the $\alpha$-th, etc.
@@
    An array is $\mathfrak{L}_{\lambda}$-indiscernible exactly when it is mutually indiscernible, i.e. each of its rows is indiscernible over the union of all other rows.
@@
    An array indexed on $I\times J$ is $\mathfrak{L}_{\mathrm{ar}}$-indiscernible exactly when each of its rows is indiscernible over the union of all other rows and, furthermore, its rows form an indiscernible sequence.
We use strongly indiscernible as a synonym for $\mathfrak{L}_{\mathrm{ar}}$-indiscernible”; note that [Ka] uses the terminology array-indiscernible instead.
@@
Let $I$ be an $\mathfrak{L}$-structure, $J,L$ be $\mathfrak{L}^{\prime}$-structures, and $E$ a set of parameters. If $c_{L}$ is based on $b_{J}$ over $a_{I}E$ and $a_{I}$ is $\mathfrak{L}$-indiscernible over $b_{J}E$, then $a_{I}$ is also $\mathfrak{L}$-indiscernible over $c_{L}E$.
@@
By naming parameters, we may assume $E=\emptyset$. Consider any finite tuples $\bar{\eta},\bar{\nu}\subseteq I$ with $\operatorname{qftp}_{\mathfrak{L}}(\bar{\eta})=\operatorname{qftp}_{\mathfrak{L}}(\bar{\nu})$, as well as a finite tuple $\bar{\ell}\subseteq L$. As $c_{L}$ is based on $b_{J}$ over $a_{I}$, there is a tuple $\bar{\jmath}\subseteq J$ with $b_{\bar{\jmath}}\equiv_{a_{I}}c_{\bar{\ell}}$. As $a_{I}$ is $\mathfrak{L}$-indiscernible over $b_{J}$, we have that $a_{\bar{\nu}}\equiv_{b_{\bar{\jmath}}}a_{\bar{\eta}}$. It follows that $a_{\bar{\nu}}\equiv_{c_{\bar{\ell}}}a_{\bar{\eta}}$, hence that $a_{\bar{\nu}}\equiv_{c_{L}}a_{\bar{\eta}}$. ∎
For an ordinal $\lambda$ and $I\subseteq\lambda$, by $\operatorname{ot}(I)$ we will mean the order type of $I$ with the order induced from $\lambda$. For $Z\subseteq\lambda\times\mu$ we write $Z_{I}\coloneqq Z\cap(I\times\mu)$ and consider it as an $\mathfrak{L}_{\operatorname{ot}(I)}$-structure in the obvious way.
By inspecting atomic $\mathfrak{L}_{\lambda}$-formulas, we easily see that if $\lambda=I\dot{\cup}J$, $\bar{\nu},\bar{\nu}^{\prime}\subseteq Z_{I}$ and $\bar{\xi},\bar{\xi}^{\prime}\subseteq Z_{J}$, then $\operatorname{qftp}_{\mathfrak{L}_{\lambda}}(\nu,\xi)=\operatorname{qftp}_{\mathfrak{L}_{\lambda}}(\nu^{\prime},\xi^{\prime})$ if and only if $\operatorname{qftp}_{\mathfrak{L}_{\operatorname{ot}(I)}}(\nu)=\operatorname{qftp}_{\mathfrak{L}_{\operatorname{ot}(I)}}(\nu^{\prime})$ and $\operatorname{qftp}_{\mathfrak{L}_{\operatorname{ot}(J)}}(\xi)=\operatorname{qftp}_{\mathfrak{L}_{\operatorname{ot}(J)}}(\xi^{\prime})$.
@@
Let $\lambda=I\dot{\cup}J$ and $Z,Z^{\prime}\subseteq\lambda\times\mu$ with $Z_{J}=Z^{\prime}_{J}$. Suppose $b_{Z^{\prime}_{I}}$ is $\mathrm{EM}_{\mathfrak{L}_{\operatorname{ot}(I)}}$-based on $a_{Z_{I}}$ over $a_{Z_{J}}E$. Let
@@
Then $c_{Z^{\prime}}$ is $\mathrm{EM}_{\mathfrak{L}_{\lambda}}$-based on $a_{Z}$ over $E$.
@@
We may assume that $E=\emptyset$. Let $\bar{\nu}\subseteq Z_{I}$, $\bar{\xi}\subseteq Z_{J}$, and $\varphi$ be such that $\varphi(x_{\bar{\nu}},x_{\bar{\xi}})\in\operatorname{EM}_{\mathfrak{L}_{\lambda}}(a_{Z})$. We have to prove that $\vDash\varphi(c_{\bar{\nu}},c_{\bar{\xi}})$.
For every $\bar{\nu}^{\prime}$ with $\operatorname{qftp}_{\mathfrak{L}_{\operatorname{ot}(I)}}(\bar{\nu}^{\prime})=\operatorname{qftp}_{\mathfrak{L}_{\operatorname{ot}(I)}}(\bar{\nu})$ we have that $\operatorname{qftp}_{\mathfrak{L}_{\lambda}}(\bar{\nu}^{\prime},\bar{\xi})=\operatorname{qftp}_{\mathfrak{L}_{\lambda}}(\bar{\nu},\bar{\xi})$, so $\vDash\varphi(a_{\bar{\nu}^{\prime}},a_{\bar{\xi}})$. Hence $\varphi(x_{\bar{\nu}^{\prime}},a_{\bar{\xi}})\in\operatorname{EM}_{\mathfrak{L}_{\operatorname{ot}(I)}}(a_{Z_{I}}/a_{Z_{J}})$ and so $\vDash\varphi(b_{\bar{\nu}},a_{\bar{\xi}})$ by the assumption on $b_{Z^{\prime}_{I}}$. This means that $\vDash\varphi(c_{\bar{\nu}},c_{\bar{\xi}})$, as required. ∎
@@
Let $\lambda=I\dot{\cup}J$ and $Z\subseteq\lambda\times\mu$. Suppose $a_{Z}$ is such that $a_{Z_{J}}$ is $\mathfrak{L}_{\operatorname{ot}(J)}$-indiscernible over $a_{Z_{I}}E$ and $a_{Z_{I}}$ is $\mathfrak{L}_{\operatorname{ot}(I)}$-indiscernible over $a_{Z_{J}}E$. Then $a_{Z}$ is $\mathfrak{L}_{\lambda}$-indiscernible over $E$.
@@
We may assume $E=\emptyset$. Let $\bar{\nu},\bar{\nu}^{\prime}\subseteq Z_{I}$ and $\bar{\xi},\bar{\xi}^{\prime}\subseteq Z_{J}$ be such that $\operatorname{qftp}_{\mathfrak{L}_{\lambda}}(\bar{\nu},\bar{\xi})=\operatorname{qftp}_{\mathfrak{L}_{\lambda}}(\bar{\nu}^{\prime},\bar{\xi}^{\prime})$. It follows easily from the assumptions that $a_{\bar{\nu},\bar{\xi}}\equiv a_{\bar{\nu},\bar{\xi}^{\prime}}\equiv a_{\bar{\nu}^{\prime},\bar{\xi}^{\prime}}$. ∎
@@
Let $A=(a_{ij})_{i<\lambda,j<\omega}$ be an array with each row consisting of compatible tuples, and $E$ a set of parameters. Then there is an array $(b_{ij})_{i<\lambda,j<\omega}$ which is $\mathfrak{L}_{\lambda}$-indiscernible over $E$ and $\mathrm{EM}_{\mathfrak{L}_{\lambda}}$-based on $A$ over $E$.